OpenStudy (anonymous):

2y to the second power -11y+5=0 solve for y

OpenStudy (anonymous):

\[2y^2-11y+5=0\] Let's try factoring (2y - 1)(y - 5)=0

OpenStudy (anonymous):

but where are you getting (2y-1) (y-5)=0? what happened to 11?

OpenStudy (anonymous):

I factored it

OpenStudy (anonymous):

ohhhhhhh got it i see because after when u foil you get that answer

OpenStudy (anonymous):

got it got it thanks